# The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxies

### v1

@@ Abstract
So far, even the highest resolution galaxy formation simulations with gravitational softening have failed to reproduce realistic life cycles of star clusters. We present the first star-by-star galaxy models of star cluster formation to account for hydrodynamics, star formation, stellar evolution and collisional gravitational interactions between stars and compact remnants using the updated sphgal+ketju code, part of the griffin-project. Gravitational dynamics in the vicinity of $>3$ M stars and their remnants are solved with a regularised integrator (ketju) without gravitational softening. Comparisons of idealised star cluster evolution with sphgal+ketju and direct N-body show broad agreement and the failure of simulations that use gravitational softening. In the hydrodynamical dwarf galaxy simulations run with sphgal+ketju, clusters up to $\sim 900$ M are formed compact (effective radii 0.1–1 pc) and their sizes increase by up to a factor of ten in agreement with previous N-body simulations and the observed sizes of exposed star clusters. The sizes increase rapidly once the clusters become exposed due to photoionising radiation. On average 63% of the gravitationally bound clusters disrupt during the first 100 Myr of evolution in the galactic tidal field. The addition of collisional dynamics reduces the fraction of supernovae in bound clusters by a factor of $\sim 2.6$, however the global star formation and outflow histories change by less than 30%. We demonstrate that the accurate treatment of gravitational encounters with massive stars enables more realistic star cluster life cycles from the earliest stages of cluster formation until disruption in simulated low-mass galaxies.
@@
pubyear: 2024pagerange: The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxiesA
@@ 1 Introduction
Young star clusters have mass functions $dN/dM\propto M^{\alpha}$ with a power-law index close to $\alpha=-2$ across different star-forming environments (Elmegreen & Falgarone, 1996; Elmegreen & Efremov, 1996; Zhang & Fall, 1999; Hunter et al., 2003; Parker & Goodwin, 2007; Fall & Chandar, 2012; Mok et al., 2019). A truncation in the mass function at high cluster masses has been suggested (Adamo et al., 2015; Messa et al., 2018; Adamo et al., 2020b), but quantifying this cut-off mass is difficult due to low-number statistics. The mass-size relation of star clusters is observed to be shallow (half-mass or half-light radii $\propto M^{\beta}$ with $\beta\lesssim 1/3$, see the data collected in Krumholz et al. 2019 and Brown & Gnedin 2021). These measurements exclude the youngest, embedded phase of early cluster evolution that is only revealed at rest-frame infrared wavelengths. The unprecedented sensitivity of JWST enables the detection of embedded star formation in the form of individual proto-stars down to sub-solar masses within the nearby galactic environments (De Marchi et al., 2017; Jones et al., 2023), thus clear improvements to the measurements of embedded cluster sizes can be expected in near future.
@@
The distribution of gas around young clusters has been used to estimate the time-scale of gas expulsion (Whitmore et al., 2011; Corbelli et al., 2017; Grasha et al., 2018; Messa et al., 2021; Hannon et al., 2022), in order to quantify the role of various stellar feedback processes in determining the initial cluster properties. Recently exposed clusters are often very young ($<5$ Myr) and have been associated with gaseous outflows (Levy et al., 2021; Sirressi et al., 2024) conceptually similar to expanding superbubbles (Watkins et al., 2023). With the exposed clusters extending to ages as low as $\sim 2$ Myr, pre-supernova feedback (HII regions and stellar winds) has been proposed as an important driver of gas expulsion in young star clusters even before the destructive supernovae (SNe) occur. Counter-examples exist as well: Kim et al. (2023) and Calzetti et al. (2023) find massive, embedded clusters that are 5–6 Myr old, indicating that massive and/or compact enough clusters can resist gas expulsion. The majority of intermediate age and old massive globular clusters (GCs) exhibit chemical variations but only in their light-element abundances (Bastian & Lardo, 2018; Gratton et al., 2019). This could be the smoking gun of gas retention in massive clusters that undergo star formation for an extended period of time while massive stars release chemically enriched stellar winds (Krause et al., 2016; Szécsi & Wünsch, 2019; Lahén et al., 2024b). What role cluster mass, compactness and metallicity play in the ability of a star cluster to continue forming stars even after the ignition of the first massive stars still remain open questions.
@@
Resolving the initial properties of star clusters and the dominant sources of feedback energy on scales of individual stars is difficult in clustered environments. High-resolution hydrodynamical simulations of galactic star formation can shed light on these complex processes that originate from small spatial scales (see e.g. Naab & Ostriker, 2017). Individual (massive) stars and their stellar feedback processes have been included in galaxy-scale simulations to study e.g. galactic interstellar medium (ISM) structure and outflows (Hu et al., 2017; Gutcke et al., 2021; Steinwandel et al., 2024b), star formation efficiency (Hislop et al., 2022) and cosmological star formation history of low-mass galaxies (Gutcke et al., 2022), clustering of stars and their feedback (Lahén et al., 2020; Smith et al., 2021; Calura et al., 2022; Gutcke, 2024; Deng et al., 2024) and the impact of runaway stars on star formation and outflows (Andersson et al., 2020; Steinwandel et al., 2023; Andersson et al., 2023). Lahén et al. (2023) introduced a numerical model (based on Hu et al., 2017 and Hu, 2019 and references therein) within the Galaxy Realizations Including Feedback From INdividual massive stars [^1] https://www.mpa-garching.mpg.de/~naab/griffin-project (griffin) project which considers every newly formed star as an individual particle. The inclusion of stellar evolution models describing the mass, energy, and metal output of massive and very massive stars throughout their lives enables the exploration of their role in the formation and evolution of star clusters star-by-star in their galactic environment. En route to building a self-consistent picture of how GCs form in a galactic context, Lahén et al. (2024b) analysed the early chemical enrichment in proto-GCs and demonstrated that massive clusters can retain gas and recycle stellar wind-material efficiently compared to SN-material.
@@
However, one key ingredient missing in previous studies of galactic-scale star formation is the collisional treatment of the individually realised stars, which we address in this paper. Gravitational two-body and multi-body interactions drive the internal evolution of star clusters through relaxation, mass-segregation, dynamical binary formation and core-collapse (Spitzer & Hart, 1971; Spitzer, 1987; Goodman & Hut, 1993), all of which are suppressed when gravitational softening is used to smooth out the gravitational potential in close encounters. Concurrent modelling of star formation, stellar feedback and collisional N-body dynamics has been successfully demonstrated in simulations of individual giant molecular clouds and star-forming regions (Wall et al., 2020; Dinnbier & Walch, 2020; Fujii et al., 2021; Grudić et al., 2021; Rieder et al., 2022).
@@
Partmann et al. (2024a) introduced recently ketju in the sphgal version of gadget-3 in a new suite of star-by-star griffin-simulations where ketju was used to compute accurate interactions of central massive black holes with surrounding individual stars while interactions between stars were still softened. Here we expand the code-implementation of Partmann et al. (2024a) by adding the ketju regularised integration regions around all forming massive stars in an entire galaxy. We execute comparison simulations with sphgal-ketju and the direct N-body code bifrost (Rantala et al., 2023, 2024). Star clusters modelled using ketju expand in size and undergo dynamical mass-loss. Hydrodynamical simulations with star formation, an evolving galactic tidal field and an accurate accounting of encounters with massive stars produce more realistic star clusters. The clusters form compact and expand rapidly once the first massive stars ignite and expel any remaining gas.
@@ 2.1 Hydrodynamical simulation code sphgal
The simulations for this study were run with the sphgal-version (Hu et al., 2014, 2016, 2017; Hu, 2019) of the gadget-3 code (Springel, 2005). We used the modern smoothed particle hydrodynamics (SPH) implementation in sphgal, which includes several improvements to the numerical accuracy of the SPH methodology as outlined e.g. in Hu et al. (2014, 2016). Recent upgrades (Lahén et al. 2023) to the code include a method for locally mass-conserving sampling of individual stars from a given stellar initial mass function (IMF) and stellar models that describe the evolution of massive stars described briefly below. Here we expand the methodology for accurate treatment of collisional gravitational dynamics introduced in Partmann et al. (2024a) to include high-accuracy ketju (Rantala et al., 2017) integration regions around all massive stars and their remnants. In the simulations, we assume an initial metallicity of $Z\sim 0.01$ $\mathrm{Z}_{\odot}$ and adopt a fiducial resolution of 4 $\mathrm{M}_{\odot}$ in gas and old stellar disk particles and $6.8\times 10^{3}$ $\mathrm{M}_{\odot}$ in dark matter. The stars formed during the simulation have masses sampled from an IMF down to the hydrogen burning limit of 0.08 $\mathrm{M}_{\odot}$.
@@ 2.1.1 Cooling, chemistry and star formation
As outlined in Hu et al. (2016) and Hu et al. (2017), cooling and chemistry of the ISM are treated in two temperature regimes. We use a chemical network at low temperatures ($<3\times 10^{4}$ K) and tabulated metal-dependent cooling rates from Wiersma et al. (2009) at high temperatures ($>3\times 10^{4}$ K). The chemistry network includes six chemical species (H2, H+, H, CO, C+, O) and free electrons, based on the methods of Nelson & Langer (1997), Glover & Mac Low (2007) and Glover & Clark (2012) closely following the implementation in the SILCC-project (Walch et al., 2015). The chemistry network takes into account the spatially resolved interstellar radiation field that is attenuated based on the gas and dust column density as outlined in Section 2.1.2. The code tracks the abundances of 13 individual elements in gas and stars: H, He, N, C, O, Si, Al, Na, Mg, Fe, S, Ca and Ne.
@@
Table 1: Description of the N-body simulations. The columns give the simulation name, the number of stars, the number of random initialisations, the total simulation time, the code used, the stellar gravitational softening length, the size of the regularised region and the initial stellar mass of particles used as centres of ketju-regions. Run hyd_KETJU_m3_e001 was limited to a total runtime of 400 Myr due to the increased computational cost.
@@ 2.1.2 Stellar feedback
Stars more massive than $9$ $\mathrm{M}_{\odot}$ are supplemented with radiation and stellar wind properties that evolve according to the stellar age. We adopt the Bonn Optimized Stellar Tracks (BoOST, Szécsi et al. 2022) that describe the evolution of stars between 9–500 $\mathrm{M}_{\odot}$. We track the far-ultraviolet and photoionising ($>13.6$ eV) luminosities, and the stellar winds characterised by the mass-loss rate, velocity and chemical composition. Photoionisation is done using the Strömgren approximation while stellar winds are released by injecting momentum and metal-enriched material into $12\times 8(\pm 2)$ gas particles in 12 equally divided solid angles around the star (see details in Hu et al. 2017; Lahén et al. 2023).
@@ 2.1.3 Regularised integrator ketju
In order to accurately model the gravitational interactions at small separations between stars in the young star clusters without gravitational softening, we utilise the ketju integration technique (Rantala et al., 2017) implemented originally in the gadget-3 code. The most recent version of ketju (Mannerkoski et al., 2023) and its regularised integrator library mstar (Rantala et al., 2020) are publicly available [^3] https://www.mv.helsinki.fi/home/phjohans/ketju as a module for the gadget-4 code (Springel et al., 2021). The gadget-3-ketju code has been used in a wide range of studies examining the dynamics of supermassive black holes in isolated mergers of early-type galaxies (Rantala et al., 2018, 2019), dark matter halos and low-mass galaxies (Partmann et al., 2024b) and gas-rich galaxies (Liao et al., 2023, 2024a, 2024b) as well as in cosmological simulations of forming early-type galaxies (Mannerkoski et al., 2021, 2022). We use the gadget-3-version of ketju as described in Mannerkoski et al. (2021) as it is directly compatible with sphgal.
@@
As the ketju-region size has to be at least $2.8$ times the gravitational softening length of the particles, we adopt here a factor of 3, i.e. $r_{\mathrm{\textsc{ketju}}}=3\epsilon_{*}$ where $\epsilon_{*}$ is the stellar softening length. ketju-regions that overlap are merged into one, thus the central regions of star cluster can host ketju-regions that are significantly larger than the single region size. The dynamics of low-mass stars below the threshold mass $m_{i}$ remains softened unless the low-mass stars are within a regularised region of a nearby massive star. The reservoir stars (before IMF-sampling is done), old disk particles, gas particles and dark matter particles are also not included in the ketju calculations. We do not use the optional star-star softening inside the ketju-regions since our N-body particles represent actual physical stars instead of macro-particles as has been the case in most previous ketju studies. For the user-given accuracy parameters we use a Gragg-Bulirsch-Stoer tolerance of $\eta_{\mathrm{GBS}}=10^{-7}$ and a end-time iteration tolerance of $\eta_{\mathrm{t}}=10^{-3}$. The simulations in this study are fully Newtonian.

### v2

@@ Abstract
So far, even the highest resolution galaxy formation simulations with gravitational softening have failed to reproduce realistic life cycles of star clusters. We present the first star-by-star galaxy models of star cluster formation to account for hydrodynamics, star formation, stellar evolution and collisional gravitational interactions between stars and compact remnants using the updated sphgal+ketju code, part of the griffin-project. Gravitational dynamics in the vicinity of $>3$ M stars and their remnants are solved with a regularised integrator (ketju) without gravitational softening. Comparisons of idealised star cluster evolution with sphgal+ketju and direct N-body show broad agreement and the failure of simulations that use gravitational softening. In the hydrodynamical simulations of idealised dwarf galaxies run with sphgal+ketju, clusters up to $\sim 900$ M form compact (effective radii 0.1–1 pc) and their sizes increase by up to a factor of ten in agreement with previous N-body simulations and the observed sizes of exposed star clusters. The sizes increase rapidly once the clusters become exposed due to photoionising radiation. On average 63% of the gravitationally bound clusters disrupt during the first 100 Myr of evolution in the galactic tidal field. The addition of collisional dynamics reduces the fraction of supernovae in bound clusters by a factor of $\sim 1.7$, however the global star formation and outflow histories change by less than 30%. We demonstrate that the accurate treatment of gravitational encounters with massive stars enables more realistic star cluster life cycles from the earliest stages of cluster formation until disruption in simulated low-mass galaxies.
@@
pubyear: 2025pagerange: The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxiesA
@@ 1 Introduction
Young star clusters have mass functions $dN/dM\propto M^{\alpha}$ with a power-law index close to $\alpha=-2$ across different star-forming environments (Elmegreen & Falgarone, 1996; Elmegreen & Efremov, 1996; Zhang & Fall, 1999; Hunter et al., 2003; Fall & Chandar, 2012; Mok et al., 2019). A truncation in the mass function at high cluster masses has been suggested (Adamo et al., 2015; Messa et al., 2018; Adamo et al., 2020b), but quantifying this cut-off mass is difficult due to low-number statistics. The mass-size relation of star clusters is observed to be shallow (half-mass or half-light radii $\propto M^{\beta}$ with $\beta\lesssim 1/3$, see the data collected in Krumholz et al. 2019 and Brown & Gnedin 2021). These measurements exclude the youngest, embedded phase of early cluster evolution that is only revealed at rest-frame infrared wavelengths. The unprecedented sensitivity of JWST enables the detection of embedded star formation in the form of individual proto-stars down to sub-solar masses within the nearby galactic environments (De Marchi et al., 2017; Jones et al., 2023), thus clear improvements to the measurements of embedded cluster sizes can be expected in near future.
@@
The distribution of gas around young clusters has been used to estimate the time-scale of gas expulsion (Whitmore et al., 2011; Corbelli et al., 2017; Grasha et al., 2018; Messa et al., 2021; Hannon et al., 2022), in order to quantify the role of various stellar feedback processes in determining the initial cluster properties. Recently exposed clusters are often very young ($<5$ Myr) and have been associated with gaseous outflows (Levy et al., 2021; Sirressi et al., 2024) conceptually similar to expanding superbubbles (Watkins et al., 2023). With the exposed clusters extending to ages as low as $\sim 2$ Myr, pre-supernova feedback (HII regions and stellar winds) has been proposed as an important driver of gas expulsion in young star clusters even before the destructive supernovae (SNe) occur. Counter-examples exist as well: Kim et al. (2023) and Calzetti et al. (2023) find massive, embedded clusters that are 5–6 Myr old, indicating that massive and/or compact enough clusters can resist gas expulsion. The majority of intermediate age and old massive globular clusters (GCs) exhibit chemical variations but only in their light-element abundances (Bastian & Lardo, 2018; Gratton et al., 2019). This could be the smoking gun of gas retention in massive clusters that undergo star formation for an extended period of time while massive stars release chemically enriched stellar winds (Krause et al., 2016; Szécsi & Wünsch, 2019; Lahén et al., 2024). What role cluster mass, compactness and metallicity play in the ability of a star cluster to continue forming stars even after the ignition of the first massive stars still remain open questions. He et al. (2019) have for instance found in molecular cloud simulations that regardless of the initial mass or the compactness of the progenitor gas clouds, they will continue forming stars for a few sound crossing times after which radiation disperses the remaining gas. However, galaxy scale simulations are needed to address the full life cycle of molecular clouds and star clusters (see e.g. the discussion in Jeffreson et al. 2024).
@@
Resolving the initial properties of star clusters and the dominant sources of feedback energy on scales of individual stars is difficult in observations of clustered environments. High-resolution hydrodynamical simulations of galactic star formation can shed light on these complex processes that originate from small spatial scales (see e.g. Naab & Ostriker, 2017). Individual (massive) stars and their stellar feedback processes have been included in galaxy-scale simulations to study e.g. galactic interstellar medium (ISM) structure and outflows (Hu et al., 2017; Steinwandel et al., 2020; Gutcke et al., 2021; Steinwandel et al., 2024b), star formation efficiency (Hislop et al., 2022), cosmological star formation history and Population III stellar feedback in low-mass galaxies (Gutcke et al., 2022; Sugimura et al., 2024; Andersson et al., 2025), clustering of stars and their feedback in idealised (Lahén et al., 2020; Smith et al., 2021; Deng et al., 2024; Lahén et al., 2024) and cosmological settings (Calura et al., 2022; Garcia et al., 2023; Gutcke, 2024; Calura et al., 2024), and the impact of runaway stars on star formation and outflows (Andersson et al., 2020; Steinwandel et al., 2023; Andersson et al., 2023). Lahén et al. (2023) introduced a numerical model (based on Hu et al., 2017 and Hu, 2019 and references therein) within the Galaxy Realizations Including Feedback From INdividual massive stars [^1] https://www.mpa-garching.mpg.de/~naab/griffin-project (griffin) project which considers every newly formed star as an individual particle. The inclusion of stellar evolution models describing the mass, energy, and metal output of massive and very massive stars throughout their lives enables the exploration of their role in the formation and evolution of star clusters star-by-star in their galactic environment. En route to building a self-consistent picture of how GCs form in a galactic context, Lahén et al. (2024) analysed the early chemical enrichment in proto-GCs and demonstrated that massive clusters can retain gas and recycle stellar wind-material efficiently compared to SN-material.
@@
However, one key ingredient missing in previous studies of galactic-scale star formation is the collisional treatment of the individually realised stars, which we address in this paper. Gravitational two-body and multi-body interactions drive the internal evolution of star clusters through relaxation, mass-segregation, dynamical binary formation and core collapse (Spitzer & Hart, 1971; Spitzer, 1987; Goodman & Hut, 1993), all of which are suppressed when gravitational softening is used to smooth out the gravitational potential in close encounters. Concurrent modelling of star formation, stellar feedback and collisional N-body dynamics has been successfully demonstrated in simulations of individual giant molecular clouds and star-forming regions (Wall et al., 2020; Dinnbier & Walch, 2020; Fujii et al., 2021; Grudić et al., 2021; Rieder et al., 2022).
@@
Partmann et al. (2025) introduced recently ketju in the sphgal version of gadget-3 in a new suite of star-by-star griffin-simulations where ketju was used to compute accurate interactions of central massive black holes with surrounding individual stars while interactions between stars were still softened. Here we expand the code-implementation of Partmann et al. (2025) by adding the ketju regularised integration regions around all forming massive stars in an entire galaxy. We execute comparison simulations with sphgal-ketju and the direct N-body code bifrost (Rantala et al., 2023, 2024). Star clusters modelled using ketju expand in size and undergo dynamical mass-loss. We then run hydrodynamical simulations with star formation and an evolving galactic tidal field resembling that of the Wolf–Lundmark–Melotte galaxy as described in Hu et al. (2017). An accurate accounting of gravitational encounters with massive stars produces more realistic star clusters. The clusters form compact and expand rapidly once the first massive stars ignite and expel any remaining gas.
@@ 2.1 Hydrodynamical simulation code sphgal
The simulations for this study were run with the sphgal-version (Hu et al., 2014, 2016, 2017; Hu, 2019) of the gadget-3 code (Springel, 2005). We used the modern smoothed particle hydrodynamics (SPH) implementation in sphgal, which includes several improvements to the numerical accuracy of the SPH methodology as outlined e.g. in Hu et al. (2014, 2016). Recent upgrades (Lahén et al. 2023) to the code include a method for locally mass-conserving sampling of individual stars from a given stellar initial mass function (IMF) and stellar models that describe the evolution of massive stars described briefly below. Here we expand the methodology for accurate treatment of collisional gravitational dynamics introduced in Partmann et al. (2025) to include high-accuracy ketju (Rantala et al., 2017) integration regions around all massive stars and their remnants. In the simulations, we assume an initial metallicity of $Z\sim 0.01$ $\mathrm{Z}_{\odot}$ and adopt a fiducial resolution of 4 $\mathrm{M}_{\odot}$ in gas and old stellar disk particles and $6.8\times 10^{3}$ $\mathrm{M}_{\odot}$ in dark matter. The stars formed during the simulation have masses sampled from an IMF down to the hydrogen burning limit of 0.08 $\mathrm{M}_{\odot}$.
@@ 2.1.1 Cooling, chemistry and star formation
As outlined in detail in Hu et al. (2016) and Hu et al. (2017), cooling and chemistry of the ISM are treated in two temperature regimes. We use a chemical network at low temperatures ($<3\times 10^{4}$ K) and tabulated metal-dependent cooling rates from Wiersma et al. (2009) at high temperatures ($>3\times 10^{4}$ K). The chemistry network includes six chemical species (H2, H+, H, CO, C+, O) and free electrons, based on the methods of Nelson & Langer (1997), Glover & Mac Low (2007) and Glover & Clark (2012) closely following the implementation in the SILCC-project (Walch et al., 2015). The chemistry network takes into account the spatially resolved interstellar radiation field that is attenuated based on the gas and dust column density as outlined in Section 2.1.2. The code tracks the abundances of 13 individual elements in gas and stars: H, He, N, C, O, Si, Al, Na, Mg, Fe, S, Ca and Ne.
@@
Table 1: Description of the N-body simulations. The columns give the simulation name, the number of stars subject to ketju-integration, the number of random initialisations, the total simulation time, the code used, the stellar gravitational softening length, the size of the regularised region and the initial stellar mass of particles used as centres of ketju-regions. Run hyd_KETJU_m3_e001 was limited to a total runtime of 400 Myr due to the increased computational cost.
@@ 2.1.2 Stellar feedback
Stars more massive than $9$ $\mathrm{M}_{\odot}$ are supplemented with radiation and stellar wind properties that evolve according to the stellar age. We adopt the Bonn Optimized Stellar Tracks (BoOST, Szécsi et al. 2022) that describe the evolution of stars between 9–500 $\mathrm{M}_{\odot}$. We track the far-ultraviolet and photoionising ($>13.6$ eV) luminosities, and the stellar winds characterised by the mass-loss rate, velocity and chemical composition. Photoionisation is implemented using the Strömgren approximation by setting the gas to be fully ionised at $10^{4}$ K within the Strömgren sphere of any photoionising star, iteratively in case overlapping HII regions occur (see Section 2.5.2 in Hu et al. 2017). Stellar winds are released by injecting momentum and metal-enriched material into $12\times 8(\pm 2)$ gas particles in 12 equally divided solid angles around the star (see details in Lahén et al. 2023).
@@ 2.1.3 Regularised integrator ketju
In order to accurately model the gravitational interactions at small separations between stars in the young star clusters without gravitational softening, we utilise the ketju integration technique (Rantala et al., 2017) implemented originally in the gadget-3 code. The most recent version of ketju (Mannerkoski et al., 2023) and its regularised integrator library mstar (Rantala et al., 2020) are publicly available [^3] https://www.mv.helsinki.fi/home/phjohans/ketju as a module for the gadget-4 code (Springel et al., 2021). The gadget-3-ketju code has been used in a wide range of studies examining the dynamics of supermassive black holes in isolated mergers of early-type galaxies (Rantala et al., 2018, 2019), dark matter halos and low-mass galaxies (Partmann et al., 2024) and gas-rich galaxies (Liao et al., 2023, 2024a, 2024b) as well as in cosmological simulations of forming early-type galaxies (Mannerkoski et al., 2021, 2022). We use the gadget-3-version of ketju as described in Mannerkoski et al. (2021) as it is directly compatible with sphgal.
@@
As the ketju-region size has to be at least $2.8$ times the gravitational softening length of the particles, we adopt here a factor of 3, i.e. $r_{\mathrm{\textsc{ketju}}}=3\epsilon_{*}$ where $\epsilon_{*}$ is the stellar softening length. ketju-regions that overlap are merged into one, thus the central regions of star cluster can host ketju-regions that are significantly larger than the single region size. The dynamics of low-mass stars below the threshold mass $m_{i}$ remains softened unless the low-mass stars are within a regularised region of a nearby massive star. The reservoir stars (before IMF-sampling is done), old disk particles, gas particles and dark matter particles are also not included in the ketju calculations. All interactions that are not between stars in a ketju-region are computed using the standard leapfrog integrator and tree force algorithm in gadget-3. We do not use the optional star-star softening inside the ketju-regions since our N-body particles represent actual physical stars instead of macro-particles as has been the case in most previous ketju studies. For the user-given accuracy parameters we use a Gragg-Bulirsch-Stoer tolerance of $\eta_{\mathrm{GBS}}=10^{-7}$ and a end-time iteration tolerance of $\eta_{\mathrm{t}}=10^{-3}$. The simulations in this study are fully Newtonian.